A proton accelerates from rest in a uniform electric field of 662 N/C. At some later time, its speed is 1.02 x 106 m/s. (a) Find the magnitude of the acceleration of the proton. m/s2 (b) How long does it take the proton to reach this speed? Hs (c) How far has it moved in that interval? (d) What is its kinetic energy at the later time?
Answer #1

a) mass of proton=1.673*10-27kg

q=1.6*1019C

so we have the formula

F=q*E=ma

so 1.673*10-27*a=1.602*10-19*662

a=6.34*1010m/s2

b) we know the formula

v=u+at

t=v/a=1.02*106/6.34*1010=1.61*10-5s

so answer is 16.1 \mus or 16\mus

c) we know the formula

v2=u2+2as

s=v2/2a=(1.02*106)2/2*6.34*1010=8.21m

so the answer is 8.21m or 8.2 m

d) we have the formula

k.e=1/2mv2=1/2*1.673*10-27*(1.02*106)2=8.70*10-16N.m

so the answer is 8.7*10-16N.m